# Classification of locality preserving symmetries on spin chains

### v1

@@ Appendix F Proof of the main Theorem
$(\delta\lambda)(g_{1},g_{2},g_{3})=h\mapsto\frac{\lambda_{\bar{g}_{1}hg_{1}}(g_{2},g_{3})\lambda_{h}(g_{1},g_{2}g_{3})}{\lambda_{h}(g_{1}g_{2},g_{3})\lambda_{h}(g_{1},g_{2})}\equiv 1.$ (G.3)
Two twisted 2-cocycles are equivalent (represent the same twisted cohomology class) if they are equal up to multiplication by a twisted 2-coboundary
$h\mapsto\frac{\epsilon_{\bar{g}_{1}hg_{1}}(g_{2})\epsilon_{h}(g_{1})}{\epsilon_{h}(g_{1}g_{2})}$ (G.4)
@@
That the phases $\omega_{j}$ form a 3-cocycle, as well as the fact that the class $[\omega_{j}]$ is independent of the choice of right restriction and the choice of fusion operators is well known, see for example [17, Appendix B] for proofs. Since a right restriction at $j^{\prime}$ can be viewed as a right restriction at $j$ with perhaps a different defect size, this also shows independence from $j$. It follows that $\operatorname{\Omega}(\alpha)$ is well defined.
@@
    If $\alpha^{\prime}$ is a decoupled symmetry then $\operatorname{\Omega}(\alpha^{\prime})=[1]$ by item 1 and $\operatorname{\Omega}(\alpha\otimes\alpha^{\prime})=\operatorname{\Omega}(\alpha)\cdot\operatorname{\Omega}{\alpha^{\prime}}=\operatorname{\Omega}(\alpha)$ by item 3. Together with the invariance of the anomaly under $\sim_{0}$, this shows that the anomaly is constant on stable equivalence classes.
@@
We first show that the class $[\lambda]\in H^{2}(G,U(1)[G])$ is independent of the choice of right restriction. Let $\alpha_{\geq}$ and $\tilde{\alpha}_{\geq}$ be right restriction of the symmetry $\alpha$. Then there are local unitaries $\{\Gamma_{g}\}_{g\in G}$ such that
$\tilde{\alpha}_{L}^{(g)}:=\mathrm{Ad}[\Gamma_{g}]\circ\alpha_{L}^{(g)}.$
@@
$\mathrm{Ad}[\tilde{\operatorname{\Psi}}_{g}(h)]=\alpha^{(h)}\circ\tilde{\alpha}_{L}^{(\bar{h}gh)}\circ\left(\alpha^{(h)}\right)^{-1}\circ\left(\tilde{\alpha}_{L}^{(g)}\right)^{-1}=\mathrm{Ad}\left[\alpha^{(h)}(\Gamma_{\bar{h}gh})\,\operatorname{\Psi}_{g}(h)\,\Gamma_{g}^{*}\right].$
@@
$\tilde{\operatorname{\Psi}}_{g}(h)=\varepsilon_{g}(h)\alpha^{(h)}\big{(}\Gamma_{\bar{h}gh}\big{)}\,\operatorname{\Psi}_{g}(h)\,\Gamma_{g}^{*}$ (I.1)
@@
Appendix J Computation of obstructions to covariant right restrictions
@@
$\alpha^{(k)}=\alpha_{\leq}^{(k)}\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[U^{(k)}_{a-1,a}]$
@@
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(g)}\circ\alpha_{\geq}^{(k)}\big{)}^{-1}\circ\mathrm{Ad}[L_{a-1}^{(k)}]^{*}$
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\left(\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}\right)$
@@
$\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}=\mathrm{Ad}[W_{a}(g,k)]$
@@
$=\mathrm{Ad}\big{[}W^{\prime}_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}(g,k)^{*}\big{]}$
@@
We can therefore take $\operatorname{\Psi}_{g}(k)=W_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}(g,k)^{*}$, which is supported on $\{a-1,a\}$.
@@
Successively applying 3-cocycle relations for elements $(k,l,\bar{l}\,\bar{k}g_{a},\bar{g}_{a}g_{a-1})$, then $(k,l,\bar{l}\,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})$, then $(k,l,\bar{l}\,\bar{k}gkl,\bar{l}\,\bar{k}\,\bar{g}g_{a})$, then $(k,\bar{k}gk,l,\bar{l}\,\bar{k}\,\bar{g}kl)$, and finally $(g,k,l,\bar{l}\,\bar{k}\,\bar{g}\,g_{a})$ this becomes
@@
Appendix K Stable equivalence is necessary
We describe a ${\mathbb{Z}}_{2}$ symmetry with trivial anomaly which is not equivalent to a decoupled symmetry through conjugation by any FDQC. This shows that stacking with decoupled symmetries is really necessary to obtain the classification result of Theorem B.1.
Consider the spin-1/2 chain ${\mathcal{A}}$ with ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{2})$ for all sites $j\in{\mathbb{Z}}$. For each site $j$ let $Z_{j}$ be the Pauli-Z operator at site $j$ and let $P^{\downarrow}_{j}=(\operatorname{\mathbbm{1}}-Z_{j})/2$. Write $CZ_{j,j^{\prime}}=\operatorname{\mathbbm{1}}-2P_{j}^{\downarrow}P_{j+1}^{\downarrow}$ for the controlled $Z$ gate acting on sites $j$ and $j^{\prime}$. Note that all these gates commute with each other.
For any finite interval $I=[a,b]$ define $U_{I}=\prod_{j=a}^{b-1}CZ_{j,j+1}$. Then $\alpha^{(-)}:=\lim_{a\uparrow\infty}\mathrm{Ad}[U_{[-a,a]}]$ defines the non-trivial component of a ${\mathbb{Z}}_{2}$-symmetry $\alpha$ with trivial anomaly.
Suppose it were possible to decouple $\alpha$ by a FDQC, then in particular there would exist a local unitary $V$ such that for all $a\in{\mathbb{N}}$ large enough we have $VU_{[-a,a]}V^{*}=U^{(L)}_{a}U^{(R)}_{a}$ with $U_{a}^{(L)}\in{\mathcal{A}}_{[-a,0]}$ and $U^{(R)}_{a}\in{\mathcal{A}}_{[1,a]}$.
@@
$\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}=2^{b}\times U_{[-a,-(b+1)]}\times CZ_{-{b+1},b+2}\times U_{[b+2,a]}$ (K.1)
for all $a>b$. We can now take $a$ and $b$ large enough so that $V\in{\mathcal{A}}_{[-b,b]}$. By invariance of the partial trace under unitary conjugation we have $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}$ is also given by Eq. (K.1). In contrast, if $VU_{[-a,a]}V^{*}=U^{(L)}_{a}U^{(R)}_{a}$ then $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,0]}\big{\{}U_{a}^{(L)}\big{\}}\times\operatorname{Tr}_{[1,b+1]}\big{\{}U_{a}^{(R)}\big{\}}$, which is incompatible with the form given by Eq. (K.1).

### v2

@@ G.2 Twisted group cohomology
$\frac{\lambda_{\bar{g}_{1}hg_{1}}(g_{2},g_{3})\lambda_{h}(g_{1},g_{2}g_{3})}{\lambda_{h}(g_{1}g_{2},g_{3})\lambda_{h}(g_{1},g_{2})}=1,\qquad\text{for all }h,g_{1},g_{2},g_{3}\in G.$ (G.6)
Two twisted 2-cocycles are equivalent (represent the same twisted cohomology class) if they are equal up to multiplication by a twisted 2-coboundary, i.e., a map
$(\delta\epsilon)_{h}(g_{1},g_{2})=\frac{\epsilon_{\bar{g}_{1}hg_{1}}(g_{2})\epsilon_{h}(g_{1})}{\epsilon_{h}(g_{1}g_{2})}$ (G.7)
@@ G.2.1 Slant product
G.2.2 Twisted group cohomology with indices in $U(1)[G^{2}]$
The $G$-module $U(1)[G^{2}]$ consists of maps $\mu:G^{2}\to U(1)$ that assign $(g,h)\mapsto\mu_{g,h}$, together with the left $G$-action
$(k\cdot\mu)_{g,h}=\mu_{\bar{k}gk,\bar{k}hk}$
and the twisted cohomology groups $H^{n}(G,U(1)[G^{2}])$ are constructed analogously as $H^{n}(G,U(1)[G^{2}])$.
Example G.3.
A 1-cocycle $\mu:G\to U(1)[G^{2}]$ is a map satisfying
$\dfrac{\mu_{\bar{k}gk,\bar{k}hk}(l)\mu_{g,h}(k)}{\mu_{g,h}(kl)}=1,\qquad g,h,k,l\in G.$ (G.9)
Similarly, a 1-coboundary is a map $\epsilon:G\to U(1)[G^{2}]$ satisfying
$\epsilon_{g,h}(k)=(\delta\nu)_{g,h}(k)=\frac{\nu_{\bar{k}gk,\bar{k}hk}}{\nu_{g,h}},$ (G.10)
for some $\nu\in U(1)[G^{2}]$.
@@ G.2.2 Twisted group cohomology with indices in $U(1)[G^{2}]$
That the phases $\omega_{j}$ form a 3-cocycle, as well as the fact that the class $[\omega_{j}]$ is independent of the choice of right restriction and the choice of fusion operators is well known, see for example [23, Appendix B] for proofs. Since a right restriction at $j^{\prime}$ can be viewed as a right restriction at $j$ with perhaps a different defect size, this also shows independence from $j$. It follows that $\operatorname{\Omega}(\alpha)$ is well defined.
@@
    If $\alpha^{\prime}$ is a decoupled symmetry then $\operatorname{\Omega}(\alpha^{\prime})=[1]$ by item 1 and $\operatorname{\Omega}(\alpha\otimes\alpha^{\prime})=\operatorname{\Omega}(\alpha)\cdot\operatorname{\Omega}(\alpha^{\prime})=\operatorname{\Omega}(\alpha)$ by item 3. Together with the invariance of the anomaly under $\sim_{0}$, this shows that the anomaly is constant on stable equivalence classes.
@@
We first show that the class $[\lambda]\in H^{2}(G,U(1)[G])$ is independent of the choice of right restriction. Let $\alpha_{\geq}$ and $\tilde{\alpha}_{\geq}$ be right restriction of the symmetry $\alpha$. Then there are local unitaries $\{W_{g}\}_{g\in G}$ such that
$\tilde{\alpha}_{\geq}^{(g)}:=\mathrm{Ad}[W_{g}]\circ\alpha_{\geq}^{(g)}.$
@@
$\mathrm{Ad}[\tilde{\operatorname{\Psi}}_{g}(h)]=\alpha^{(h)}\circ\tilde{\alpha}_{\geq}^{(\bar{h}gh)}\circ\left(\alpha^{(h)}\right)^{-1}\circ\left(\tilde{\alpha}_{\geq}^{(g)}\right)^{-1}=\mathrm{Ad}\left[\alpha^{(h)}(W_{\bar{h}gh})\,\operatorname{\Psi}_{g}(h)\,W_{g}^{*}\right].$
@@
$\tilde{\operatorname{\Psi}}_{g}(h)=\varepsilon_{g}(h)\alpha^{(h)}\big{(}W_{\bar{h}gh}\big{)}\,\operatorname{\Psi}_{g}(h)\,W_{g}^{*}$ (I.1)
@@
Appendix J Proof of Proposition E.7
Independence of the phases of the fusion operators. Suppose $\alpha_{\geq}$ is a covariant right restriction of $\alpha$ with fusion operators $\Phi$ leading to
$\alpha^{(k)}\big{(}\Phi(\bar{k}gk,\bar{k}hk)\big{)}=\mu_{g,h}(k)\times\Phi(g,h).$
The fusion operators are determined only up to phase, so we can use $\widetilde{\Phi}(g,h)=\xi(g,h)\Phi(g,h)$ instead, leading to
$\alpha^{(k)}\big{(}\widetilde{\Phi}(\bar{k}gk,\bar{k}hk)\big{)}=\tilde{\mu}_{g,h}(k)\times\widetilde{\Phi}(g,h)$
with
$\tilde{\mu}_{k}(g,h)=\frac{\xi_{\bar{k}gk,\bar{k}hk}}{\xi_{g,h}}\mu_{k}(g,h).$
Hence $\mu$ and $\tilde{\mu}$ differ by a twisted 1-coboundary (G.10), thus $[\mu]=[\tilde{\mu}]\in\mathfrak{K}$.
Independence of right restriction. Let $\alpha_{\geq}$ and $\beta_{\geq}$ be covariant right restrictions of $\alpha$. Then there are unitaries $U(g)$ such that $\beta_{\geq}^{(g)}=\mathrm{Ad}[U(g)]\circ\alpha_{\geq}^{(g)}$ and it follows from covariance that there are phases $c_{g}(k)$ such that $\alpha^{(k)}\big{(}U(\bar{k}gk)\big{)}=c_{g}(k)\times U(g).$ By computing $\alpha^{(kl)}\big{(}U(\overline{kl}gkl)\big{)}$ in two different ways we find moreover that the $c_{g}(k)$ satisfy the twisted 1-cocycle law $c_{g}(kl)=c_{g}(k)c_{\bar{k}gk}(l).$ By choosing different phases $\widetilde{U}(g)=\eta(g)U(g)$ we get new
$\tilde{c}_{g}(k)=\frac{\eta_{\bar{k}gk}}{\eta_{g}}c_{g}(k).$
Suppose $\alpha_{\geq}$ has fusion operators $\Phi$ leading to $\alpha^{(k)}\big{(}\Phi(\bar{k}gk,\bar{k}hk)\big{)}=\mu_{g,h}(k)\times\Phi(g,h).$ Then $\beta_{\geq}$ has fusion operators
$\widetilde{\Phi}(g,h)=U(g)\,\alpha_{\geq}^{(g)}\big{(}U(h)\big{)}\,\Phi(g,h)\,U(gh)^{*}$
and we compute (using covariance of $\alpha_{\geq}$)
$\alpha^{(k)}\big{(}\widetilde{\Phi}(\bar{k}gk,\bar{k}hk)\big{)}$ $=\alpha^{(k)}\big{(}U(\bar{k}gk)\big{)}\,\alpha^{(g)}_{\geq}\big{(}\alpha^{(k)}\big{(}U(\bar{k}hk)\big{)}\big{)}\,\alpha^{(k)}\big{(}\Phi(\bar{k}gk,\bar{k}hk)\big{)}\,\alpha^{(k)}\big{(}U(\bar{k}ghk)^{*}\big{)}$
$=\frac{c_{g}(k)c_{h}(k)}{c_{gh}(k)}\mu_{g,h}(k)\times U(g)\,\alpha^{(g)}_{\geq}\big{(}U(h)\big{)}\,\Phi(g,h)\,U(gh)^{*}$
$=\frac{c_{g}(k)c_{h}(k)}{c_{gh}(k)}\mu_{g,h}(k)\times\widetilde{\Phi}(g,h).$
This shows that
$\tilde{\mu}_{g,h}(k)=\frac{c_{g}(k)c_{h}(k)}{c_{gh}(k)}\mu_{g,h}(k)$
for some representative $c$ of a class $[c]\in H^{1}(G,U(1)[G])$. This shows $[\tilde{\mu}]=[\mu]\in\mathfrak{K}$.
The proof of items 1 to 4 are similar to the corresponding proofs of items 1 through 4 in Appendix H.
Appendix K Computation of obstructions to covariant right restrictions
@@
$\alpha^{(k)}=\alpha_{\leq}^{(k)}\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[V^{(k)}_{a-1,a}]$
@@
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[V_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[V_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(g)}\circ\alpha_{\geq}^{(k)}\big{)}^{-1}\circ\mathrm{Ad}[L_{a-1}^{(k)}]^{*}$
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\left(\mathrm{Ad}[V_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[V_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}\right)$
@@
$\mathrm{Ad}[V_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[V_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}=\mathrm{Ad}[W_{a}(g,k)]$
@@
$=\mathrm{Ad}\big{[}W^{\prime}_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}_{a}(g,k)^{*}\big{]}$
@@
We can therefore take $\operatorname{\Psi}_{g}(k)=W^{\prime}_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}_{a}(g,k)^{*}$, which is supported on $\{a-1,a\}$.
@@
Successively applying 3-cocycle relations for elements $(k,l,\bar{l}\,\bar{k}g_{a},\bar{g}_{a}g_{a-1})$, then $(k,l,\bar{l}\,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})$, then $(k,l,\bar{l}\,\bar{k}gkl,\bar{l}\,\bar{k}\,\bar{g}g_{a})$, then $(k,\bar{k}gk,l,\bar{l}\,\bar{k}\,\bar{g}g_{a})$, and finally $(g,k,l,\bar{l}\,\bar{k}\,\bar{g}\,g_{a})$ this becomes
@@
Appendix L Stable equivalence is necessary
We describe a ${\mathbb{Z}}_{2}$ symmetry with trivial anomaly which is not equivalent to a decoupled symmetry [37]. This shows Theorem B.1 would not hold true if we had replaced stable equivalence" by "equivalence".
Consider the spin chain ${\mathcal{A}}$ with ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{2})$ for all sites $j\in{\mathbb{Z}}$. For each site $j$ let $Z_{j}$ be the $Z$-Pauli matrix, i.e. $\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ acting at site $j$ and let $P^{\downarrow}_{j}=(\operatorname{\mathbbm{1}}-Z_{j})/2$. Write $CZ_{j,j+1}=\operatorname{\mathbbm{1}}-2P_{j}^{\downarrow}P_{j+1}^{\downarrow}$ for the controlled $Z$ gate acting on sites $j$ and $j+1$. Note that all these gates commute with each other.
For any finite interval $I=[a,b]$ define $U_{I}=\prod_{j=a}^{b-1}CZ_{j,j+1}$. Then $\alpha^{(-1)}:=\lim_{a\uparrow\infty}\mathrm{Ad}[U_{[-a,a]}]$ defines the non-trivial component of a ${\mathbb{Z}}_{2}$-symmetry $\alpha$ with trivial anomaly.
Suppose it were possible to decouple $\alpha$ by a FDQC, then in particular there would exist a local unitary $V$ such that for all $a\in{\mathbb{N}}$ large enough we have $VU_{[-a,a]}V^{*}=U^{L}_{a}U^{R}_{a}$ with $U_{a}^{L}\in{\mathcal{A}}_{[-a,0]}$ and $U^{R}_{a}\in{\mathcal{A}}_{[1,a]}$.
@@
$\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}=2^{b}\times U_{[-a,-(b+1)]}\times CZ_{-{b+1},b+2}\times U_{[b+2,a]}$ (L.1)
for all $a>b$. We can now take $a$ and $b$ large enough so that $V\in{\mathcal{A}}_{[-b,b]}$. By invariance of the partial trace under unitary conjugation we have $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}$ is also given by Eq. (L.1). In contrast, if $VU_{[-a,a]}V^{*}=U^{L}_{a}U^{R}_{a}$ then $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,0]}\big{\{}U_{a}^{L}\big{\}}\times\operatorname{Tr}_{[1,b+1]}\big{\{}U_{a}^{R}\big{\}}$, which is incompatible with the form given by Eq. (L.1).